OpenStudy (anonymous):

Find S22 for -6 + (-4) + (-2) + 0 …

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

@Eyad

OpenStudy (anonymous):

You try out now. 1st term A = -6, find the difference, and find T22 . Hence find S22. try up once.

OpenStudy (anonymous):

A=-6 ,d =2 S22 = a + (n-1)d =-6 + ( 21)2 = 36 Sn = n/2( a+l) Then , S22 = 22/2 ( -6 +36) =330

OpenStudy (anonymous):

Thx bruh you help me with my exam @Eyad

OpenStudy (anonymous):

@luc

OpenStudy (anonymous):

Yw :3 ,Good Luck